# MTL-KD: Multi-Task Learning Via Knowledge Distillation for Generalizable Neural Vehicle Routing Solver

### v2

@@ 6.1 Main Results
HGS-PyVRP CVRP 15.53(*) 62.07(*) 119.54(*) VRPTW 24.35(*) 90.61(*) 166.47(*)
@@
HGS-PyVRP VRPL 15.58(*) 63.55(*) 122.68(*) VRPB - - -
OR-Tools 16.00(2.69%) 67.53(6.26%) 128.16(4.46%) 11.97(*) 47.76(*) 88.57(*)
MT-POMO(S.T.) 16.29(4.52%) 71.46(12.66%) 149.5521.90(%) 12.47(4.21%) 50.08(4.85%) 99.56(12.41%)
MT-POMO(M+aug8) 15.85(1.67%) 68.95(8.49%) 138.97(13.27%) 12.04(0.63%) 48.49(1.52%) 95.35(7.64%)
MVMoE(S.T.) 16.36(4.94%) 78.79(23.97%) 192.29(56.74%) 12.42(3.80%) 71.99(50.75%) 186.49(110.54%)
MVMoE(M+aug8) 15.81(1.45%) 74.67(17.50%) 177.89(45.00%) 12.01(0.31%) 66.26(38.73%) 167.26(88.84%)
MTL-KD96 16.14(3.58%) 65.45(2.99%) 126.11(2.79%) 12.39(3.51%) 46.12(-3.44%) 86.90(-1.89%)
MTL-KD128 16.12(3.45%) 65.48(3.03%) 126.66(3.25%) 12.38(3.41%) 45.99(-3.70%) 86.99(-1.78%)
MTL-KD(R3C200)128 15.82(1.50%) 64.5246 (1.53%) 124.59(1.56%) 12.02(0.43%) 44.58(-6.66%) 83.66(-5.55%)
HGS-PyVRP OVRP 9.71(*) 35.30(*) 66.10(*) OVRPTW 13.95(*) 48.15(*) 82.98(*)
OR-Tools 9.84(1.38%) 37.81(7.11%) 70.38(6.48%) 14.38(3.06%) 52.48(9.00%) 90.03(8.49%)
MT-POMO(S.T.) 10.66(9.83%) 44.62(26.41%) 92.78(40.36%) 15.56(11.53%) 71.63(48.76%) 147.84(78.16%)
MT-POMO(M+aug8) 10.17(4.74%) 41.93(18.78%) 85.27(29.00%) 14.85(6.43%) 66.74(38.61%) 135.99(63.87%)
MVMoE(S.T.) 10.77(10.94%) 49.35(39.81%) 135.30(104.70%) 15.70(12.54%) 82.81(71.99%) 215.21(159.34%)
MVMoE(M+aug8) 10.14(4.42%) 46.23(30.97%) 116.82(76.74%) 14.78(5.90%) 76.28(58.43%) 195.64(135.76%)
MTL-KD96 10.41(7.21%) 38.73(9.72%) 72.73(10.03%) 15.11(8.32%) 53.74(11.61%) 94.12(13.42%)
MTL-KD128 10.46(7.19%) 38.89(10.17%) 72.70(9.98%) 15.08 (8.07%) 53.89(11.91%) 93.96(13.22%)
MTL-KD(R3C200)128 10.05(3.53%) 37.7934 (7.07%) 71.40(8.03%) 14.53 (4.12%) 52.08(8.15%) 92.40(11.35%)
@@
We evaluat the performance of our model on the training tasks, and the results are shown in Table 1. The results demonstrate that our proposed MTL-KD model exhibits excellent performance on the seen training tasks, particularly showcasing a more pronounced advantage when dealing with larger-scale problems. This validates the effectiveness of our proposed knowledge distillation training framework and model architecture.
@@
HGS-PyVRP OVRPL 9.67(*) 34.70(*) 65.38(*) VRPLTW 24.44(*) 91.86(*) 174.79(*)
@@
HGS-PyVRP OVRPLTW 14.00(*) 47.97 (*) 83.68(*) OVRPB - - -
@@
OR-Tools VRPBL 12.02(*) 47.93 (*) 89.82 (*) VRPBTW 25.41(*) 97.77(*) 194.69(*)
@@
OR-Tools OVRPBL 8.35 (*) 29.60(*) 54.30 (*) OVRPBTW 14.38(*) 51.88 (*) 90.86 (*)
@@
OR-Tools VRPBLTW 25.34 (*) 103.17(*) 189.31(*) OVRPBLTW 14.25(*) 52.37(*) 92.16(*)
@@
We also evaluate the performance of our model on real-world datasets, including Set-X (medium and large scale) from CVRPLIB[25] for CVRP and the Solomon dataset[26] for VRPTW. We compare our model against several single-task and multi-task models, and the experimental results are presented in Tables 5, 6, and 7, AppendixC. Our model consistently outperforms other single-task and multi-task models across all three real-world datasets, demonstrating particularly strong performance on the large-scale Set-X dataset, where the gap is significantly smaller than other baselines. This indicates the strong applicability of our approach to real-world scenarios.
@@ 6.2 Ablation Study
While heavy decoders are generally considered to have significant potential for large-scale generalization, our model’s learning teacher model, POMO, primarily excels at its specific training scale and lacks cross-scale generalization capabilities. This naturally raises a crucial question: will the student model also be limited to learning knowledge specific to a particular scale, thereby restricting its generalization ability? To investigate this, we compare the performance of several single-task teacher models with our MTL-KD model on the same dataset, as shown in Figure 4. In contrast to the teacher models, the student model demonstrates a notable capacity for scale generalization. This observation indicates that the MTL-KD model, during the learning process, does not simply replicate the teacher’s knowledge but rather effectively adapts to variations in scale, consequently achieving robust generalization performance.
@@
To highlight the effectiveness of our knowledge distillation approach, we compared the MTL-KD model with a LEHD model trained using RL. Given the high computational cost associated with training the LEHD model with RL, we were only able to train it at a scale of 20. Subsequently, we evaluated the average gap of both models on all seen and unseen tasks, with the experimental results presented in Table 3. The results indicate that the RL method, due to its limitation to training on small scales, struggles to leverage the full potential of the LEHD model, leading to significantly poor performance. In contrast, our distillation-based MTL-KD model demonstrates a clear performance advantage.
@@
Training task RLMT-LEHD 21.7834% 27.2457% 38.2808% 50.0606%
@@
unseen task RLMT-LEHD 31.2513% 37.1404% 48.9074% 62.2226%
@@
To validate the effectiveness of our proposed R3C method, we conduct experiments on CVRP, VRPL, OVRP, and VRPTW at a scale of 100. We perform an ablation study on the random reordering of the subtour external order component, with the following comparative experiments: random sampling on the original solution only (RS); random reordering of the subtour external order followed by sampling (RS+Ro); and for CVRP and VRPL, we also analyze the impact of adding a random flipping operation (F) since it does not affect the legality of the solution. The experimental results are shown in Figure 5. Incorporating the random reordering operation significantly improves the iterative performance during reconstruction, which benefits from the ability of random reordering to enhance the diversity of sampled subproblems. Furthermore, the random flipping operation can moderately increase the diversity of sampled solutions when only initial solution random sampling is performed; however, its impact is minimal with random reordering, further confirming the effectiveness of the random reordering operation.
@@ 7 Conclusion
This paper implements a heavy decoder model in the multi-task VRP domain and achieves label-free training of this heavy decoder model through a multi-task knowledge distillation method. The proposed approach demonstrates excellent performance on seen tasks, unseen tasks, and real-world datasets, exhibiting significant large-scale generalization ability, thereby validating the effectiveness of our method. Furthermore, the proposed R3C method further enhances the model’s performance. However, the structure of the heavy decoder leads to high computational complexity, which is a limitation. Future work will explore how to implement more efficient and robust model architectures in the multi-task domain.
@@ 20. Guidelines:
  23. Answer: [No]
  24. Justification: We will release the dataset and code upon manuscript acceptance, with all experimental details documented in the paper to ensure full reproducibility of our results.
@@ 48. Answer: [Yes]
  49. Justification: This work focuses on solving and optimizing combinatorial optimization (CO) problems, which holds the potential to significantly enhance real-world applications. It appears to have no negative social impact. More discussion refers to Appendix E.
@@ 58. Answer: [Yes]
  59. Justification: We list the licenses in Appendix D.
@@ Appendix A Related work
A.1 Autoregressive Neural Solvers
Autoregressive neural solvers have emerged as a prominent research direction in the field of NCO. These methods typically employ an encoder-decoder architecture, where the encoder generates node embeddings, and the decoder dynamically produces a sequence of nodes.
@@ A.1 Autoregressive Neural Solvers
Early works were based on Pointer Networks[27], gradually evolving from supervised learning to reinforcement learning training[28, 29]. Subsequently, the Transformer framework was introduced and has become a mainstream paradigm[7]. POMO[8] further leverages the symmetry of the VRP, optimizing the baseline calculation in RL training and enhancing the exploration diversity of the model. Subsequent research has largely focused on improving the Transformer framework [30, 31, 10, 9, 32, 33, 34]. However, these approaches predominantly adopt a heavy encoder light decoder structure, which poses challenges in handling large-scale problems. To enhance the capability of addressing large instances, some methods continue to build upon HELD. Examples include ELG[35], which integrates local search strategies; INVIT[36], which utilizes K-nearest neighbor multi-view embeddings to enhance local information; and ICAM[37], which introduces instance-conditional adaptation to better perceive problem scales. On the other hand, some studies[15, 16] have explored heavy decoder architectures. By dynamically capturing the relationships between remaining nodes during the decoding process, these methods demonstrate superior generalization performance on large-scale problems. However, their high computational demands, often necessitating training on labeled data, represent a significant limitation.Furthermore, decomposition-based methods[38, 39, 40, 17] warrant attention. These approaches decompose large-scale problems into multiple smaller subproblems for solving, and autoregressive solvers can be employed at the lower level to address these subproblems.
A.2 Multi-Task Learning for Solving VRPs
@@
To enhance the generalizability of neural solvers across diverse VRP variants, several existing methods have proposed unified multi-task models. For instance, Wang et al.[41] employs a multi-armed bandit approach to achieve efficient training across multiple combinatorial optimization problems. Lin et al.[42] pre-train a Transformer backbone on the Traveling Salesman Problem and fine-tune it for specific VRP variants, thereby extending its applicability to a subset of VRP problems. These methods represent initial attempts at cross-problem learning but are limited in scope, focusing primarily on a small number of problems. Building on these foundations, MT-POMO [11] views VRP variants as combinations of distinct attributes, enabling unified representation and zero-shot generalization across multiple VRP problems, achieving notable performance on ten tasks. MVMoE [12] further boosts model capacity and multi-task performance by introducing a mixture-of-experts model. RouteFinder[13] enhances multi-task training efficiency and performance through mixed batch training and enables rapid adaptation to unseen tasks via efficient adapter layers. CaDA [43] improves the model’s ability to handle diverse tasks within a multi-task framework by incorporating constraint awareness and a dual-branch structure. However, these studies have predominantly focused on small-scale problems, lacking extensive exploration of large-scale generalization capabilities.
@@ Appendix B Problem Definition and Settings
Our VRP variant problem settings follow the work of Zhou et al. [12]. Detailed settings are as follows:
  • Coordinates: All node coordinates are uniformly sampled from the 2D space $[0,1)\times[0,1)$.
@@
Appendix C Performance on Real-World Instances
Our test results on Real-World Instances datasets are presented in Tables 5, 6, and 7. All models are trained on instances of size $N=100$. During inference, data augmentation is applied, and a greedy rollout strategy is adopted by default. Some results for comparative models are sourced from MVMoE [12] and RouteFinder [13].
Table 5: Results on small-scale CVRPLIB instances (Set-X) [25]. All models are trained on instances of size $N=100$, and inference utilized data augmentation[8] and greedy rollout by default.

### v3

@@ 6.1 Main Results
HGS-PyVRP
CVRP
15.53(*) 62.07(*) 119.54(*)
VRPTW
24.35(*) 90.61(*) 166.47(*)
@@
HGS-PyVRP
VRPL
15.58(*) 63.55(*) 122.68(*)
OVRP
9.71(*) 35.30(*) 66.10(*)
OR-Tools 16.00(2.69%) 67.53(6.26%) 128.16(4.46%) 9.84(1.38%) 37.81(7.11%) 70.38(6.48%)
MT-POMO(S.T.) 16.29(4.52%) 71.46(12.66%) 149.5521.90(%) 10.66(9.83%) 44.62(26.41%) 92.78(40.36%)
MT-POMO(M+aug8) 15.85(1.67%) 68.95(8.49%) 138.97(13.27%) 10.17(4.74%) 41.93(18.78%) 85.27(29.00%)
MVMoE(S.T.) 16.36(4.94%) 78.79(23.97%) 192.29(56.74%) 10.77(10.94%) 49.35(39.81%) 135.30(104.70%)
MVMoE(M+aug8) 15.81(1.45%) 74.67(17.50%) 177.89(45.00%) 10.14(4.42%) 46.23(30.97%) 116.82(76.74%)
RF-Transformer(M+aug8) 15.88(1.93%) 68.59(7.93%) 135.15(10.17%) 10.11(4.11%) 42.88(21.48%) 84.40(27.68%)
RF-MTPOMO(M+aug8) 15.93(2.23%) 68.33(7.52%) 134.53(9.66%) 10.17(4.70%) 41.60(17.86%) 82.14(24.26%)
RF-MVMoE(M+aug8) 15.90(2.06%) 68.39(7.62%) 138.60(12.98%) 10.13(4.29%) 41.18(16.65%) 81.82(23.78%)
CaDa(M+aug8) 15.89(2.00%) 176.04(177.01%) 536.71(337.49%) 10.10(4.04%) 187.72(431.78%) 442.57(569.55%)
MTL-KD96 16.14(3.58%) 65.45(2.99%) 126.11(2.79%) 10.41(7.21%) 38.73(9.72%) 72.73(10.03%)
MTL-KD128 16.12(3.45%) 65.48(3.03%) 126.66(3.25%) 10.46(7.19%) 38.89(10.17%) 72.70(9.98%)
MTL-KD(R3C200)128 15.82(1.50%) 64.5246 (1.53%) 124.59(1.56%) 10.05(3.53%) 37.7934 (7.07%) 71.40(8.03%)
HGS-PyVRP
VRPB
OVRPTW
13.95(*) 48.15(*) 82.98(*)
OR-Tools 11.97(*) 47.76(*) 88.57(*) 14.38(3.06%) 52.48(9.00%) 90.03(8.49%)
MT-POMO(S.T.) 12.47(4.21%) 50.08(4.85%) 99.56(12.41%) 15.56(11.53%) 71.63(48.76%) 147.84(78.16%)
MT-POMO(M+aug8) 12.04(0.63%) 48.49(1.52%) 95.35(7.64%) 14.85(6.43%) 66.74(38.61%) 135.99(63.87%)
MVMoE(S.T.) 12.42(3.80%) 71.99(50.75%) 186.49(110.54%) 15.70(12.54%) 82.81(71.99%) 215.21(159.34%)
MVMoE(M+aug8) 12.01(0.31%) 66.26(38.73%) 167.26(88.84%) 14.78(5.90%) 76.28(58.43%) 195.64(135.76%)
MTL-KD96 12.39(3.51%) 46.12(-3.44%) 86.90(-1.89%) 15.11(8.32%) 53.74(11.61%) 94.12(13.42%)
MTL-KD128 12.38(3.41%) 45.99(-3.70%) 86.99(-1.78%) 15.08 (8.07%) 53.89(11.91%) 93.96(13.22%)
MTL-KD(R3C200)128 12.02(0.43%) 44.58(-6.66%) 83.66(-5.55%) 14.53 (4.12%) 52.08(8.15%) 92.40(11.35%)
@@
We evaluate the performance of our model on the training tasks, and the results are shown in Table 1. The results demonstrate that our proposed MTL-KD model exhibits excellent performance on the seen training tasks, particularly showcasing a more pronounced advantage when dealing with larger-scale problems. This validates the effectiveness of our proposed knowledge distillation training framework and model architecture.
@@
HGS-PyVRP
OVRPL
9.67(*) 34.70(*) 65.38(*)
VRPLTW
24.44(*) 91.86(*) 174.79(*)
@@
HGS-PyVRP
OVRPLTW
14.00(*) 47.97 (*) 83.68(*)
OVRPB
@@
OR-Tools
VRPBL
12.02(*) 47.93 (*) 89.82 (*)
VRPBTW
25.41(*) 97.77(*) 194.69(*)
@@
OR-Tools
OVRPBL
8.35 (*) 29.60(*) 54.30 (*)
OVRPBTW
14.38(*) 51.88 (*) 90.86 (*)
@@
OR-Tools
VRPBLTW
25.34 (*) 103.17(*) 189.31(*)
OVRPBLTW
14.25(*) 52.37(*) 92.16(*)
@@
We also evaluate the performance of our model on real-world datasets, including Set-X (medium and large scale) from CVRPLIB [38] for CVRP and the Solomon dataset [39] for VRPTW. We compare our model against several single-task and multi-task models, and the experimental results are presented in Tables 6, 7, and 8, AppendixD. Our model consistently outperforms other single-task and multi-task models across all three real-world datasets, demonstrating particularly strong performance on the large-scale Set-X dataset, where the gap is significantly smaller than other baselines. This indicates the strong applicability of our approach to real-world scenarios.
@@ 6.2 Ablation Study
While heavy decoders are generally considered to have significant potential for large-scale generalization, the teacher model, POMO, primarily excels at its specific training scale and lacks cross-scale generalization capabilities. This naturally raises a crucial question: Will the student model also be limited to learning knowledge specific to a particular scale, thereby restricting its generalization ability? To investigate this, we compare the performance of several single-task teacher models with our MTL-KD model on the same dataset, as shown in Figure 4. In contrast to the teacher models, the student model demonstrates a notable capacity for scale generalization. This observation indicates that the MTL-KD model, during the learning process, does not simply replicate the teacher’s knowledge but rather effectively adapts to scale variations, consequently achieving robust generalization performance.
@@
To highlight the effectiveness of our knowledge distillation approach, we train the proposed multi-task LEHD model under two distinct training paradigms: knowledge distillation (MTL-KD) and reinforcement learning (MTL-RL). Given the high computational cost associated with training the LEHD model with RL, we are only able to train it at a scale of 20. Subsequently, we evaluate the average gap of both models in all tasks seen and unseen , with the experimental results presented in Table 3. The results indicate that the RL method, due to its limitation to training on small scales, struggles to leverage the full potential of the LEHD model, leading to significantly poorer performance. In contrast, our distillation-based MTL-KD model demonstrates a clear performance advantage.
@@
Training task MTL-RL 21.7834% 27.2457% 38.2808% 50.0606%
@@
unseen task MTL-RL 31.2513% 37.1404% 48.9074% 62.2226%
@@
To validate the effectiveness of our proposed R3C method, we conduct experiments on CVRP, VRPL, OVRP, and VRPTW at scale 100. We perform an ablation study on the random reordering of the subtour external order component, with the following comparative experiments: random sampling on the original solution only (RS); random reordering of the subtour external order followed by sampling (RS+Ro); and for CVRP and VRPL, we also analyze the impact of adding a random flipping operation (F) since it does not affect the legality of the solution. The experimental results are shown in Figure 5. Incorporating the random reordering operation significantly improves the iterative performance during reconstruction, which benefits from the ability of random reordering to enhance the diversity of sampled subproblems. Furthermore, the random flipping operation can moderately increase the diversity of sampled solutions when only initial solution random sampling is performed; however, its impact is minimal with random reordering, further confirming the effectiveness of the random reordering operation.
@@ 7 Conclusion
This paper presents a heavy decoder model for the multi-task VRP domain and achieves label-free training of this model through a multi-task knowledge distillation method. The proposed approach has demonstrated excellent performance on both seen and unseen tasks, as well as real-world datasets, exhibiting significant large-scale generalization ability and thereby validating the effectiveness of our method. Furthermore, the proposed R3C method has further enhanced the model’s performance. However, the structure of the heavy decoder results in high computational complexity. Future work will explore how to implement more efficient and robust model architectures in the multi-task domain.
Acknowledgments and Disclosure of Funding
This work was supported in part by the National Natural Science Foundation of China (Grant 62476118, Grant 72271168), the Natural Science Foundation of Guangdong Province (Grant 2024A1515011759), the Guangdong Science and Technology Program (Grant 2024B1212010002), the Guangdong Basic and Applied Basic Research Foundation (Grant 2024A1515012485), and the Center for Computational Science and Engineering at Southern University of Science and Technology.
@@ 20. Guidelines:
  23. Answer: [Yes]
  24. Justification: See the abstract for the link to the code and trained models.
@@ 48. Answer: [Yes]
  49. Justification: This work focuses on solving and optimizing combinatorial optimization (CO) problems, which holds the potential to significantly enhance real-world applications. It appears to have no negative social impact. More discussion refers to Appendix F.
@@ 58. Answer: [Yes]
  59. Justification: We list the licenses in Appendix E.
@@ Appendix A Related work
A.1 Single-task Autoregressive Neural Solver
@@
Research on high-performance autoregressive neural solvers for VRP is a hot topic in NCO. It is typically based on the encoder-decoder structure, where the encoder generates node embeddings, and the decoder is used to dynamically generate node sequences. Early works are based on Pointer Networks [40], gradually evolving from supervised learning to reinforcement learning training [41, 42]. Subsequently, the Transformer framework is introduced and becomes a mainstream paradigm [10]. POMO [11] further leverages the symmetry of the VRP, optimizing the baseline calculation in RL training and enhancing the exploration diversity of the model. Subsequent research largely focuses on improving the Transformer framework [43, 44, 13, 12, 45, 46, 47]. However, these approaches predominantly adopt the Heavy Encoder and Light Decoder (HELD) structure, which poses challenges in handling large-scale problems. To enhance the capability of solving large instances, some methods continue to build upon HELD. Examples include ELG [48], which integrates local search strategies; INVIT [49], which utilizes K-nearest neighbor multi-view embeddings to enhance local information; and ICAM [50], which introduces instance-conditional adaptation to better perceive problem scales. On the other hand, some studies [27, 28] explore heavy decoder architectures. By dynamically capturing the relationships between remaining nodes during the decoding process, these methods demonstrate superior generalization performance on large-scale problems. However, their high computational demands, often necessitating training on labeled data, represent a significant limitation. Furthermore, some methods [51, 52, 53, 29] decompose large-scale problems into multiple smaller subproblems for solving, and autoregressive neural solvers can be employed at the lower level to address these subproblems.
A.2 Multi-Task Autoregressive Neural Solver
@@
To enhance the generalizability of neural solvers across diverse VRP variants, several existing methods have proposed unified multi-task models. For example, Wang and Yu [54] employ a multi-armed bandit approach to achieve efficient training across multiple combinatorial optimization problems. Lin et al.[55] pre-train a Transformer backbone on the Traveling Salesman Problem and fine-tune it for specific VRP variants, thereby extending its applicability to a subset of VRP problems. These methods represent initial attempts at cross-problem learning but are limited in scope, focusing primarily on a small number of problems. Building on these foundations, MT-POMO [23] views VRP variants as combinations of distinct attributes, enabling unified representation and zero-shot generalization across multiple VRP variants, achieving notable performance on ten tasks. MVMoE [24] further boosts model capacity and multi-task performance by introducing a mixture-of-experts model. RouteFinder [25] enhances multi-task training efficiency and performance through mixed batch training and enables rapid adaptation to unseen tasks via efficient adapter layers. CaDA [37] improves the model’s ability to handle diverse tasks within a multi-task framework by incorporating constraint awareness and a dual-branch structure. However, these studies have predominantly focused on small-scale problems, lacking extensive exploration of large-scale generalization capabilities.
@@ Appendix B Problem Definition and Settings
Our VRP variant problem settings follow the work of Zhou et al. [24]. The detailed settings are as follows:
  • Coordinates: All node coordinates are uniformly sampled from the 2D space $0,1)\times[0,1)$.
@@
Appendix C Sensitivity Analysis on R3C Re-optimization Segment Length
To confirm the efficacy of our random segment length selection in the R3C strategy, we compare it against fixed segment lengths ($k\in\{10,20,30,40,50\}$) on 100-node instances for four VRP variants (CVRP, OVRP, VRPL, VRPTW). As shown in Table 5, the random length consistently yields the best performance across all problems. We attribute this advantage to the increased solution diversity provided by varying segment lengths, which effectively prevents the search from getting trapped in local optima, a common issue with fixed segment sampling.
Table 5: Results after 50 iterations of R3C with different Re-optimization Segment Lengths on 100-node instances.
Problem segment=10 segment=20 segment=30 segment=40 segment=50 Random
CVRP 16.01 15.96 15.94 15.91 15.91 15.88
OVRP 10.34 10.25 10.21 10.17 10.20 10.17
VRPL 16.09 16.03 16.00 15.97 15.98 15.90
VRPTW 26.02 25.86 25.77 25.65 25.69 25.55
Appendix D Performance on Real-World Instances
Our test results on real-world instance datasets are presented in Tables 6, 7, and 8. All models are trained on instances of size $N=100$. During inference, data augmentation is applied, and a greedy rollout strategy is adopted by default. Some results for comparative models are sourced from MVMoE [24] and RouteFinder [25].
Table 6: Results on small-scale CVRPLIB instances (Set-X) [38]. All models are trained on instances of size $N=100$, and inference utilizes data augmentation [11] and greedy rollout by default.